OpenStudy (anonymous):

square root of 10 and 6 divided by square root of 3

OpenStudy (anonymous):

\[\sqrt{10*6} / \sqrt{3} or \sqrt{10+6}/\sqrt{3}\]

OpenStudy (anonymous):

@Purplerainbowcherry its square root of 10 and square root of 6 its two radicals at the top then divided by square root of 3

OpenStudy (anonymous):

\[\frac{ \sqrt{10}+\sqrt{6} }{ \sqrt{3} }\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay i got the answer

OpenStudy (anonymous):

\[\sqrt{10} = \sqrt{2} * \sqrt{5}\] \[\sqrt{6} = \sqrt{2} * \sqrt{3}\] so basically : \[\frac{ \sqrt{2}*\sqrt{5}+\sqrt{2}*\sqrt{3} }{ \sqrt{3} } \] and as such you can take \[\sqrt{2}\] out since it is is common : \[\frac{ \sqrt{2}(\sqrt{5}+\sqrt{3}) }{ \sqrt{3} } \] next, you multiply by \[\sqrt{3} \] in the denominator and the numerator in order to rationalise so : \[\frac{ \sqrt{2}(\sqrt{5}+\sqrt{3}) }{ \sqrt{3} } * \frac{ \sqrt{3} }{ \sqrt{3} }\]

OpenStudy (anonymous):

for the denominators , you get square root of 9 which is 3, so your denominator becomes 3

OpenStudy (anonymous):

for the numerators, you simply multiply and simplify and get : \[\sqrt{30} + 3\sqrt{2} \]